Science Report – February 13th

Science Report 13 February 2019

Crew 205 – International Emerging Space Leaders

Submitted and prepared by GreenHab Officer Nathan Hadland

1.) ISRU Study: Geological characterization of regolith samples began with observations under the optical and dissection microscopes. We have also been performing spectral analysis using a spectrometer to get the initial elemental characterization. Analysis of pH and redox potential of samples to determine the suitability of the substrate to host plant growth will be investigated in the coming days. Additionally, the Arabidopsis thaliana seeds were prepped using a washing protocol and plated on an agar germination plate. We are waiting on germination to occur. Dead plant material from the GreenHab was taken and dried using an oven in the Science Dome and subsequently weighed for the cellulosic ethanol ISRU study.

2.) GreenHab: Dead and live samples of cucumbers were taken from the GreenHab and wet mounts were prepared and viewed under the optical microscope. We found evidence of degraded cell structure, likely from the fungi growing on the dead plants. The live tissue comparatively had all of the expected components of plant cells including stoma and clear cell walls. The carrots that were not harvested were moved from around the grow bed and consolidated into one section of the planter.

3.) Astronomy: The variable star SY MON was observed on 11 Feb 2019. The observation session was one 60 s exposure using the V filter. On 12 Feb 2019, the magnitude value was measured for SY MON using AstroImage J software. The magnitude value was found to be 14.147. The value will be submitted to the AAVSO. Another observation session was submitted on 12 Feb 2019. The session was for the ASASSN-V J081823.00-111138.9 variable star with the aim of drawing a light curve. The observation session was 9 exposures in the V filter and 9 exposures in the B filter. The V filter exposures were 60 seconds, and the B filters exposures were 120 seconds. However, due to the bad weather, no images were captured.

4.) EVA Project: The primary research technique was changed after testing analysis methods from the mission plan during EVA’s of the first 2 days. A new research focus has been adopted and is currently under development. In practice, a “score” will be assigned to each EVA to determine its effectiveness by evaluating various parameters. More data will be available in upcoming reports.

5.) Effective Leadership: Role rotations will begin tomorrow and the first survey conducted on the current commander, N. Larrea will be collected tonight.

6.) Developing Guidelines (SOPs): Water conservation procedures, improved communications procedures, and EVA preparation procedures are being developed. An inventory of the Science Dome, GreenHab, and the RAM are also being compiled as well as an improved list of food provided to crews.

Copyright © The Mars Society. All rights reserved. | Main Site